You decided to build a dog house run in back corner your yard. If you only have 60 feet of fencing, what is the maximum area that you can enclose? What dimensions produce the maximum area

Answer by Alan3354(69443)

You decided to build a dog house run in back corner your yard. If you only have 60 feet of fencing, what is the maximum area that you can enclose? What dimensions produce the maximum area
-------------
Max area is a circle.
r = 60/(2pi)
Area = pi*r^2 = 900/pi sq feet
=~ 286.48 sq ft